# Aggregation Models with Optimal Weights for Distributed Gaussian Processes

### v1

@@ 2.2.1 Distributed GP training
Distibuted GPs (Ng and Deisenroth,, 2014; Deisenroth and Ng,, 2015; Liu et al.,, 2018; Yin et al.,, 2020) enable the parallel training of GPs across different experts to harness greater computing power and accelerate the training process for large-scale datasets. Assuming the same model as in Section 2.1.1, where $f(\cdot)\sim\mathcal{GP}(0,k(\cdot,\cdot^{\prime}))$ and $y_{i}=f(\mathbf{x}_{i})+\epsilon_{i}$ with the i.i.d. noise $\epsilon_{i}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$. We first partition the global training dataset $\mathcal{D}:=\{\mathbf{X},\mathbf{y}\}$ of size $n$ into $M$ local datasets $\mathcal{D}_{i}=\{\mathbf{X}_{i},\mathbf{y}_{i}\}$, each of size $n_{i}$ for $i=1,\ldots,M$. Each local dataset $\mathcal{D}_{i}$ corresponds to an expert $\mathcal{M}_{i}$. Clearly, $n=\sum_{i=1}^{M}n_{i}$. We suppose that all local datasets share the same hyperparameters $\bm{\theta}$, as defined in Section 2.1.2. The global objective for distribued GP training is to maximize the global marginal log-likelihood $\log p(\mathbf{y}|\mathbf{f},\bm{\theta})$, where $\mathbf{f}=f(\mathbf{X})$. Assuming independence among all the experts $\{\mathcal{M}_{i}\}_{i=1}^{M}$, the global marginal log-likelihood is approximated as follows
@@
The factorized approximation of the log-likilihood in eq. 13 degenerates the full covariance matrix $\mathbf{K}_{\mathbf{f}\mathbf{f}}=k(\mathbf{X},\mathbf{X})$ into a diagonal block matrix $\text{diag}[\mathbf{K}_{\mathbf{f}_{1},\mathbf{f}_{1}},\ldots,\mathbf{K}_{\mathbf{f}_{M},\mathbf{f}_{M}}]$, reducing the time complexity to $\mathcal{O}(n_{i}^{3})$. This factorized training method (Ng and Deisenroth,, 2014; Deisenroth and Ng,, 2015), termed as FACT, is one common approach to distributed GP training. In addition to FACT, various optimization methods have been proposed. Federated Averaging (FedAvg) (McMahan et al.,, 2017) is a popular aggregation strategy that trains a global model by averaging updates from multiple local datasets, FedProx (Li et al.,, 2020) generalizes FedAvg by adding a proximal term to address the heterogeneity in federated networks. The Alternating Direction of Multipliers Method (ADMM) (Boyd et al.,, 2011) reformulates the optimization problem as a nonconvex consensus problem with newly introduced local hyperparameters and the global hyperparameter. Proximal ADMM (pxADMM) (Hong et al.,, 2016) aims to reduce communication overhead and computational time, and Xie et al., (2019) derive a closed-form solution for distributed GP training based on pxADMM.
@@ 2.2.2 Aggregated prediction
Common aggregation methods for GP experts include product-of-experts (PoE) (Hinton,, 2002), generalised product-of-experts (gPoE) (Cao and Fleet,, 2014), Bayesian committee machine (BCM) (Tresp,, 2000), robust Bayesian committee machine (rBCM) (Deisenroth and Ng,, 2015), generalized robust Bayesian committee machine (grBCM) (Liu et al.,, 2018), and nested pointwise aggregation of experts (NPAE) (Rullière et al.,, 2018). The joint mean $\mu_{\mathcal{A}}$ and joint precision $\sigma_{\mathcal{A}}^{-2}$ for PoE, gPoE, BCM, rBCM are given by
@@
where $\mu_{i}(\cdot)$ and $\sigma_{i}^{-2}(\cdot)$ represent the local mean and local precision of expert $\mathcal{M}_{i}$, while $\sigma_{**}^{-2}(\cdot)=k(\cdot,\cdot)+\sigma_{\epsilon}^{2}$ is the prior variance, serving as a correction term for the BCM family. The weights, $\beta_{i}$, are defined as follows: $\beta_{i}=1$ for PoE and BCM, $\beta_{i}=1/M$ for gPoE, and $\beta_{i}=0.5(\log\sigma_{**}^{2}(\cdot)-\log\sigma_{i}^{2}(\cdot))$ for rBCM.
@@
NPAE leverages the covariances between experts to ensure consistent predictions. Consequently, the joint mean and joint variance are computed as follows:
@@
with
$\mathrm{cov}[\mu_{i}(\cdot),y_{*}(\cdot)]=\mathbf{K}_{(\cdot),i}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{K}_{\mathbf{f}_{i}(\cdot)},$ (17a)
$\mathrm{cov}[\mu_{i}(\cdot),\mu_{j}(\cdot)]=\begin{cases}\mathbf{K}_{(\cdot)\mathbf{f}_{i}}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{j}}\widetilde{\mathbf{K}}_{\mathbf{f}_{j}\mathbf{f}_{j}}^{-1}\mathbf{K}_{\mathbf{f}_{j}(\cdot)},&i\neq j,\\ \mathbf{K}_{(\cdot),i}\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}^{-1}\mathbf{K}_{\mathbf{f}_{i}(\cdot)},&i=j,\end{cases}$ (17b)
where $\mu_{i}(\cdot)$ represents the local mean of expert $\mathcal{M}_{i}$, $\mathbf{K}_{(\cdot)\mathbf{f}_{i}}=k(\cdot,\mathbf{X}_{i})=\mathbf{K}_{\mathbf{f}_{i}(\cdot)}^{\top}$, $\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{j}}=k(\mathbf{X}_{i},\mathbf{X}_{j})$, $\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}=\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{i}}+\sigma_{\epsilon}^{2}\mathbf{I}_{n_{i}}$. $\mathbf{K}_{\mathcal{A}(\cdot)}=\big{[}\mathrm{cov}[\mu_{i}(\cdot),y_{*}(\cdot)]\big{]}_{i=1}^{M}=\mathbf{K}_{(\cdot)\mathcal{A}}^{\top}\in\mathbb{R}^{M}$, $\mathbf{K}_{\mathcal{A}\mathcal{A}}=\big{[}\mathrm{cov}[\mu_{i}(\cdot),\mu_{j}(\cdot)]\big{]}_{i,j=1}^{M}\in\mathbb{R}^{M\times M}$, $\bm{\mu}_{\mathcal{A}}=\big{[}\mu_{1}(\cdot),\ldots,\mu_{M}(\cdot)\big{]}^{\top}\in\mathbb{R}^{M}$.
@@ 3 GP regression with OptiCom
In this section, we detail the algorithms for GP regression using the optimized combination technique. First, we introduce the combination technique and the optimized combination technique in Section 3.1 and Section 3.2, respectively. Then, we present the algorithm for applying OptiCom to GP in Section 3.3.
@@ 3.1 Combination technique (CT)
Combination technique (CT) (Griebel et al.,, 1990; Hegland et al.,, 2007), first introduced in (Smolyak,, 1963), is an efficient tool for approximating the sparse grid spaces. When partial projection operators commute, as in interpolation with tensor product spaces, the combination technique provides the exact sparse grid solution (Griebel et al.,, 1990). Sparse grids of level $\eta$ and dimension $d$ are defined as (Garcke,, 2013)
@@
Here, $\underline{l}=(l_{1},\ldots,l_{d})$, $|\underline{l}|_{1}=\sum_{j=1}^{d}l_{j}$, where $l_{j}\in\mathbb{N}^{+}$ for all $j=1,\ldots d$. $\Omega_{l_{1}}^{1}\times\cdots\times\Omega_{l_{d}}^{1}=\{(\omega_{l_{1}},\ldots,\omega_{l_{d}})|\omega_{l_{j}}\in\Omega_{l_{j}}^{1}\;\text{for}\;j\in\{1,\ldots,d\}\}$ denotes the n-ary Cartesian product over $d$ one-dimensional set $\Omega_{l_{j}}^{1}\subset\mathbb{R}$, $j=1,\ldots,d$. We suppose $\Omega_{l_{j}}^{1}$ consists of $2^{l_{j}-1}$ uniformly distributed points over a fixed interval, i.e., $\Omega_{l_{j}}^{1}$ over the interval $[0,1]$ is given by
$\Omega_{l_{j}}^{1}=\left\{\frac{i}{2^{l_{j}}}:i=1,\ldots,2^{l_{j}-1}\right\}.$ (19)
Each set of grids $\Omega_{\underline{l}}\subset\mathbb{R}^{d}$ is associated with a piecewise $d$-linear basis function $\bm{\phi}_{\underline{l},\underline{h}}(\cdot)$, which is defined as the product of the one-dimensional basis functions:
$\bm{\phi}_{\underline{l},\underline{h}}(\mathbf{x}):=\prod_{j=1}^{d}\phi_{l_{j},h_{j}}(x_{j}),\quad h_{j}=1,\ldots,2^{l_{j}-1}$ (20)
where $\mathbf{x}=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$ is a $d$-dimensional point, $\underline{h}=(h_{1},\ldots,h_{d})$, $\phi_{l_{j},h_{j}}(\cdot)$ is a piecewise linear hierarchical basis shown in Figure 7(a) in Appendix and defined as follows:
$\phi_{l_{j},h_{j}}(x)=\begin{cases}1-\lvert 2^{l_{j}}x-h_{j}\rvert,&x\in\big{[}\frac{h_{j}-1}{2^{l_{j}}},\frac{h_{j}+1}{2^{l_{j}}}\big{]},\\ 0,&\text{otherwise}.\end{cases}$ (21)
These basis functions can define function spaces $V_{\underline{l}}$ on the grids $\Omega_{\underline{l}}$,
$V_{\underline{l}}:=\text{span}\{\bm{\phi}_{\underline{l},\underline{h}}:h_{j}=1,\ldots,2^{l_{j}}-1,j=1,\ldots,d\}.$ (22)
Then the hierarchical increment spaces $W_{\underline{l}}$ can be defined by
@@
Therefore, the function spaces can be represented by the hierarchical increment spaces
$V_{\underline{l}}=\bigoplus_{\underline{k}\;\leq\;\underline{l}}W_{\underline{k}},$ (25)
and each function $f_{\underline{l}}\in V_{\underline{l}}$ can be uniquely represented by
$f_{\underline{l}}(\mathbf{x})=\sum_{|\underline{l}|_{1}\leq\eta+d-1}\sum_{\underline{h}\in\mathrm{B}_{\underline{l}}}\alpha_{\underline{k},\underline{h}}\bm{\phi}_{\underline{k},\underline{h}}(\mathbf{x})$ (26)
with hierarchical coefficients $\alpha_{\underline{l},\underline{h}}\in\mathbb{R}$. The combination technique linearly combines the discrete solutions $f_{\underline{l}}(\mathbf{x})$ from the partial grids $\Omega_{\underline{l}}$ according to the combination formula (see Figure 7(b) in Appendix for an example):
${\color[rgb]{0,0,1}{f_{\eta}^{c}(\mathbf{x})=\sum_{\eta\leq|\underline{l}|_{1}\leq\eta+d-1}(-1)^{\eta+d-1-|\underline{l}|_{1}}\binom{d-1}{|\underline{l}|_{1}-\eta}f_{\underline{l}}(\mathbf{x}),}}$ (27)
where the function $f_{\eta}^{c}$ is in the sparse gris space $V_{\eta}^{s}$ of level $\eta$ defined by
$V_{\eta}^{s}:=\bigoplus_{|\underline{l}|_{1}\leq\eta+d-1}W_{\underline{l}}.$ (28)
@@ 3.2 Optimized combination technique (OptiCom)
Optimized Combination Technique (OptiCom) (Hegland et al.,, 2002; Garcke,, 2006; Griebel et al.,, 2015) selects the “best possible” combination coefficients adaptively so that an optimal combination approximation is obtained. More specifically, we aim to minimize the functional
@@
${\color[rgb]{0,0,1}{f_{\eta}^{c}(\mathbf{x})=\sum_{\eta\leq|\underline{l}|_{1}\leq\eta+d-1}c_{\underline{l}}f_{\underline{l}}(\mathbf{x}).}}$ (32)
For the regression problem (Garcke,, 2006), we are seeking the solution of the optimization problem $f^{*}=\arg\min_{f\in V}R(f)$ with

### v2

@@ 2.2 Distributed GP
While SVGP reduces complexity from $\mathcal{O}(n^{3})$ to $\mathcal{O}(nm^{2})$, this can still be impractical when $n$ is extremely large such as billions of observations. Distributed learning provides a natural solution by partitioning the data across $M$ computational units (experts) and processing them in parallel, further reducing per-unit complexity to $O(nm^{2}/M)$.
This section introduces the distributed GP framework that motivates our main contributions. Section 2.2.1 describes how local experts are trained independently on partitioned local datasets. Section 2.2.2 reviews existing aggregation methods for distributed GPs and highlights their strengths and limitations.
@@ 2.2.1 Distributed GP training
Distributed GPs (Deisenroth and Ng, 2015) enable the parallel training of GPs across different experts to accelerate the computing process for large-scale datasets. Note that distributed GPs are not inherently private because they involve data sharing or centralized aggregation, whereas federated GPs preserve privacy by keeping raw data local and only exchanging model updates. Assuming the same model as in Section 2.1.1, we first partition the global training dataset $\mathcal{D}:=\{\mathbf{X},\mathbf{y}\}$ of size $n$ into $M$ local datasets $\mathcal{D}_{i}=\{\mathbf{X}_{i},\mathbf{y}_{i}\}$, each of size $n_{i}$ for $i=1,\ldots,M$. Each local dataset $\mathcal{D}_{i}$ corresponds to an expert $\mathcal{M}_{i}$. Clearly, $n=\sum_{i=1}^{M}n_{i}$. We suppose that all local datasets share the same hyperparameters $\bm{\theta}$. The global objective for distributed GP training is to maximize the global marginal log-likelihood $\log p(\mathbf{y}|\mathbf{f},\bm{\theta})$, where $\mathbf{f}=f(\mathbf{X})$. Assuming independence among all the experts $\{\mathcal{M}_{i}\}_{i=1}^{M}$, the global marginal log-likelihood is approximated as:
@@
The factorized approximation of the log-likelihood in eq. 3 (FACT) approximates the full covariance matrix $\mathbf{K}_{\mathbf{f}\mathbf{f}}=k(\mathbf{X},\mathbf{X})$ by a diagonal block matrix $\text{diag}[\mathbf{K}_{\mathbf{f}_{1},\mathbf{f}_{1}},\ldots,\mathbf{K}_{\mathbf{f}_{M},\mathbf{f}_{M}}]$, reducing the time complexity to $\mathcal{O}(n_{i}^{3})$. Other optimization techniques include Federated Averaging (FedAvg) (McMahan et al., 2017), which averages local model updates; FedProx (Li et al., 2020), which improves FedAvg by addressing client heterogeneity; ADMM (Boyd et al., 2011), which solves a consensus optimization problem by introducing auxiliary variables and dual updates.
@@ 2.2.2 Aggregated prediction
After training local experts, we aggregate local predictions to form a global prediction. Let $\mu_{i}(\cdot)$ and $\sigma_{i}^{2}(\cdot)$ denote the predictive mean and variance from expert $\mathcal{M}_{i}$. Common aggregation methods for GP experts include product-of-experts (PoE) (Hinton, 2002), generalised product-of-experts (gPoE) (Cao and Fleet, 2014), Bayesian committee machine (BCM) (Tresp, 2000), robust Bayesian committee machine (rBCM) (Deisenroth and Ng, 2015), generalized robust Bayesian committee machine (grBCM) (Liu et al., 2018), and nested pointwise aggregation of experts (NPAE) (Rullière et al., 2018). PoE, gPoE, BCM and rBCM produce a global prediction using a weighted average of local precisions, the joint mean $\mu_{\mathcal{A}}$ and joint precision $\sigma_{\mathcal{A}}^{-2}$ are given by
@@
where $\sigma_{**}^{-2}(\cdot)=k(\cdot,\cdot)+\sigma_{\epsilon}^{2}$ is the prior variance, serving as a correction term for the BCM family. The weights, $\beta_{i}$, are defined as follows: $\beta_{i}=1$ for PoE and BCM, and $\beta_{i}=0.5(\log\sigma_{**}^{2}(\cdot)-\log\sigma_{i}^{2}(\cdot))$ for gPoE and rBCM.
@@
NPAE leverages the covariances between experts to ensure consistent predictions. The joint mean and joint variance are computed as follows:
@@
where $\mathbf{K}_{(\cdot)\mathbf{f}_{i}}=k(\cdot,\mathbf{X}_{i})=\mathbf{K}_{\mathbf{f}_{i}(\cdot)}^{\top}$, $\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{j}}=k(\mathbf{X}_{i},\mathbf{X}_{j})$, $\widetilde{\mathbf{K}}_{\mathbf{f}_{i}\mathbf{f}_{i}}=\mathbf{K}_{\mathbf{f}_{i}\mathbf{f}_{i}}+\sigma_{\epsilon}^{2}\mathbf{I}_{n_{i}}$. $\mathbf{K}_{\mathcal{A}(\cdot)}=[\mathrm{cov}[\mu_{i}(\cdot),y_{*}(\cdot)]]_{i=1}^{M}$ == $\mathbf{K}_{(\cdot)\mathcal{A}}^{\top}\in\mathbb{R}^{M}$, $\mathbf{K}_{\mathcal{A}\mathcal{A}}=\big[\mathrm{cov}[\mu_{i}(\cdot),\mu_{j}(\cdot)]\big]_{i,j=1}^{M}\in\mathbb{R}^{M\times M}$, $\bm{\mu}_{\mathcal{A}}=\big[\mu_{1}(\cdot),\ldots,\mu_{M}(\cdot)\big]^{\top}\in\mathbb{R}^{M}$.
Among these methods, PoE, gPoE, BCM, and rBCM are computationally efficient but assume independence between experts, leading to inconsistent predictions as the dataset size grows. The grBCM method addresses consistency by introducing a global expert but requires training augmented models $\mathcal{M}_{i}^{+}$ for $i=2,\ldots,M$, each on datasets of size $n_{c}+n_{i}$, significantly increasing computational cost. NPAE accounts for correlations through the full covariance matrix $\mathbf{K}_{\mathcal{A}\mathcal{A}}$ in eq. 6, achieving consistency but at $O(M^{3})$ cost for inverting this matrix, which becomes prohibitive for large $M$.
These limitations motivate our approach: we seek an aggregation method that (1) incorporates expert correlations for consistency, (2) maintains computational efficiency comparable to simple baselines, and (3) naturally applies to both exact GP and SVGP settings. Our key insight introduced in Section 4 is to leverage the optimized combination technique (OptiCom) from sparse grid theory, which is introduced in Section 3, to derive optimal aggregation weights that address these requirements.
@@ 3 GP regression with OptiCom
This section introduces GP regression using the optimized combination technique (OptiCom). We first review the standard combination technique (CT) in Section 3.1, followed by OptiCom in Section 3.2, and then show how it is applied to GP regression.
@@ 3.1 Combination technique (CT)
Combination technique (CT) (Griebel et al., 1990; Hegland et al., 2007), first introduced in (Smolyak, 1963), is an efficient tool for approximating the sparse grid spaces. When partial projection operators commute, as in interpolation with tensor product spaces, the combination technique provides the exact sparse grid solution. Sparse grids of level $\eta$ and dimension $d$ are defined as (Garcke, 2013)
@@
Here, $\underline{l}=(l_{1},\ldots,l_{d})$, $|\underline{l}|_{1}=\sum_{j=1}^{d}l_{j}$, where $l_{j}\in\mathbb{N}^{+}$ for all $j=1,\ldots d$. $\Omega_{l_{1}}^{1}\times\cdots\times\Omega_{l_{d}}^{1}=\{(\omega_{l_{1}},\ldots,\omega_{l_{d}})|\omega_{l_{j}}\in\Omega_{l_{j}}^{1}\;\text{for}\;j\in\{1,\ldots,d\}\}$ denotes the n-ary Cartesian product over $d$ one-dimensional set $\Omega_{l_{j}}^{1}\subset\mathbb{R}$, $j=1,\ldots,d$. We suppose $\Omega_{l_{j}}^{1}$ consists of $2^{l_{j}-1}$ uniformly distributed points over a fixed interval, i.e., $\Omega_{l_{j}}^{1}$ over the interval $[0,1]$ is given by $\Omega_{l_{j}}^{1}=\left\{\frac{i}{2^{l_{j}}}:i=1,\ldots,2^{l_{j}-1}\right\}$.
Each set of grids $\Omega_{\underline{l}}\subset\mathbb{R}^{d}$ is associated with a piecewise $d$-linear basis function $\bm{\phi}_{\underline{l},\underline{h}}(\cdot)$ defined as:
$\bm{\phi}_{\underline{l},\underline{h}}(\mathbf{x}):=\prod_{j=1}^{d}\phi_{l_{j},h_{j}}(x_{j}),\quad h_{j}=1,\ldots,2^{l_{j}-1},\quad\phi_{l_{j},h_{j}}(x)=\begin{cases}1-\lvert 2^{l_{j}}x-h_{j}\rvert,&x\in\big[\frac{h_{j}-1}{2^{l_{j}}},\frac{h_{j}+1}{2^{l_{j}}}\big],\\ 0,&\text{otherwise},\end{cases}$ (8)
where $\mathbf{x}=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$ is a $d$-dimensional point, $\underline{h}=(h_{1},\ldots,h_{d})$, $\phi_{l_{j},h_{j}}(\cdot)$ is a piecewise linear hierarchical basis shown in Figure 7(a).
These basis functions can define function spaces $V_{\underline{l}}:=\text{span}\{\bm{\phi}_{\underline{l},\underline{h}}:h_{j}=1,\ldots,2^{l_{j}}-1,j=1,\ldots,d\}$ on the grids $\Omega_{\underline{l}}$, then the hierarchical increment spaces $W_{\underline{l}}$ can be defined by
@@
Therefore, the function spaces can be represented by the hierarchical increment spaces $V_{\underline{l}}=\bigoplus_{\underline{k}\;\leq\;\underline{l}}W_{\underline{k}}$ and each function $f_{\underline{l}}\in V_{\underline{l}}$ can be uniquely represented by $f_{\underline{l}}(\mathbf{x})=\sum_{|\underline{l}|_{1}\leq\eta+d-1}\sum_{\underline{h}\in\mathrm{B}_{\underline{l}}}\alpha_{\underline{l},\underline{h}}\bm{\phi}_{\underline{l},\underline{h}}(\mathbf{x})$ with hierarchical coefficients $\alpha_{\underline{l},\underline{h}}\in\mathbb{R}$. The space $\oplus_{i=1}^{n}S_{i}$ is the direct sum of the spaces $S_{1},\ldots,S_{n}$ if and only if all nonzero vectors $s_{i}\in S_{i}$ (for $i=1,\ldots,n$) are linearly independent, and $\underline{k}\leq\underline{l}$ indicates that each component of $\underline{k}$ is less than or equal to the corresponding component of $\underline{l}$. The combination technique linearly combines the discrete solutions $f_{\underline{l}}(\mathbf{x})$ from the partial grids $\Omega_{\underline{l}}$ according to the combination formula (see Figure 7):
$f_{\eta}^{c}(\mathbf{x})=\sum_{\eta\leq|\underline{l}|_{1}\leq\eta+d-1}(-1)^{\eta+d-1-|\underline{l}|_{1}}\binom{d-1}{|\underline{l}|_{1}-\eta}f_{\underline{l}}(\mathbf{x}),$ (11)
where the function $f_{\eta}^{c}$ is in the sparse grid space $V_{\eta}^{s}:=\bigoplus_{|\underline{l}|_{1}\leq\eta+d-1}W_{\underline{l}}$.
@@ 3.2 Optimized combination technique (OptiCom)
Optimized Combination Technique (OptiCom) (Hegland, 2002; Garcke, 2006) selects the “best possible” combination coefficients adaptively so that an optimal combination approximation is obtained. More specifically, we aim to minimize the functional
@@
$f_{\eta}^{c}(\mathbf{x})=\sum_{\eta\leq|\underline{l}|_{1}\leq\eta+d-1}c_{\underline{l}}f_{\underline{l}}(\mathbf{x}).$ (15)
For the regression problem (Garcke, 2006), we are seeking the solution of the optimization problem $f^{*}=\arg\min_{f\in V}R(f)$ with